How many grams of mannitol does the patient receive in 10 hours?
1
Expert's answer
2020-12-01T09:15:09-0500

A patient receives 100 mL of 20 % (m/v) mannitol solution every hour. How many grams of mannitol does the patient receive in 10 hours?

Proportion 1:

100 mL – 1 hr

x mL – 10 hr

"x = \frac{100 \times 10}{1} = 1000 \;mL"

Proportion 2:

100 mL – 20 g

1000 mL – y

"y = \frac{20 \times 1000}{100} = 200 \;g"

Answer: the patient receive 200 g of mannitol in 10 hours.